# Carleman Estimates and Controllability of Forward Stochastic Parabolic Equations with General Dynamic Boundary Conditions

### v1

@@ 1 Introduction and Main Results
The surface diffusion term $\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\Gamma})$ on the boundary in (1.7) plays a crucial role in deriving the Carleman estimate (1.6), specifically in the proof of Lemma 2.1—Step 2 in Appendix A. This term is essential for absorbing the following problematic integral from the right-hand side:
$\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}\psi||\nabla_{\Gamma}z_{\Gamma}|^{2}\,d\sigma dt.$
However, the study of the controllability of (1.7) with the following dynamic boundary condition:
$\begin{cases}dy_{\Gamma}+\partial_{\nu}^{\mathcal{A}}y\,dt=(a_{2}y_{\Gamma}+B_{2}\cdot\nabla_{\Gamma}y_{\Gamma})\,dt+v_{2}\,dW(t)&\textnormal{on}\;\Sigma,\\ y_{\Gamma}=y|_{\Gamma}&\textnormal{on}\;\Sigma,\end{cases}$
remains an open problem. In contrast, it is well known that for $N=1$, the controllability of stochastic parabolic equations can be studied under the above type of boundary condition; see [5].
@@
In [31], controllability results are provided for semilinear stochastic parabolic equations with Dirichlet boundary conditions and globally Lipschitz semilinear terms. Extending these results to equations with dynamic boundary conditions, where the semilinear terms involve gradient terms in both the bulk and the boundary, is a natural next step. However, for a broad class of nonlinearities, as discussed in [12, 13], this remains an open problem due to the lack of compactness embedding in stochastic partial differential equations.
Remark 1.6.
[…]
$\|z(0,\cdot)\|^{2}_{L^{2}_{\mathcal{F}_{0}}(\Omega;L^{2}(G))}+\|z_{\Gamma}(0,\cdot)\|^{2}_{L^{2}_{\mathcal{F}_{0}}(\Omega;L^{2}(\Gamma))}\leq C\,\mathbb{E}\iint_{Q_{0}}z^{2}\,dxdt.$
Alternatively, one may establish the unique continuation property:
$z=0\quad\textnormal{in}\;Q_{0},\,\;\textnormal{a.s.}\quad\Longrightarrow\quad(z_{T},z_{\Gamma,T})=(0,0)\quad\textnormal{in}\;G\times\Gamma,\,\;\textnormal{a.s.}$
For partial results related to the above problem, we refer the reader to [8, 24]. Introducing the additional controls $v_{1}$ and $v_{2}$ in the diffusion terms of equation (1.7), as discussed in Section 3, allows us to prove only a weaker observability inequality (the inequality (3.8)) and a weak unique continuation property (the property (3.7)) for the adjoint system (3.1).
In this paper, we use the following lemma to perform integration by parts in space (see, e.g., [11, 19]).
@@
For any $F\in L^{2}(G;\mathbb{R}^{N})$ and $F_{\Gamma}\in L^{2}(\Gamma;\mathbb{R}^{N})$, we define the extension of the operator divergence of $F$ (resp., tangential divergence of $F_{\Gamma}$) as the continuous linear operator on $H^{1}(G)$ (resp., on $H^{1}(\Gamma)$) as follows:
$\nabla\cdot F:H^{1}(G)\longrightarrow\mathbb{R},\quad\textbf{v}\longmapsto-\int_{G}F\cdot\nabla\textbf{v}\,dx+\langle F\cdot\nu,\textbf{v}|_{\Gamma}\rangle_{H^{-1/2}(\Gamma),H^{1/2}(\Gamma)},$
$\nabla_{\Gamma}\cdot F_{\Gamma}:H^{1}(\Gamma)\longrightarrow\mathbb{R},\quad\textbf{v}_{\Gamma}\longmapsto-\int_{\Gamma}F_{\Gamma}\cdot\nabla_{\Gamma}\textbf{v}_{\Gamma}\,d\sigma.$
Moreover, we have the following estimates:
$\left|\langle\nabla\cdot F,\textbf{v}\rangle_{(H^{1}(G))^{\prime},H^{1}(G)}\right|\leq C_{1}\|F\|_{L^{2}(G;\mathbb{R}^{N})}\|\textbf{v}\|_{H^{1}(G)},\quad\text{for all }\;F\in L^{2}(G;\mathbb{R}^{N}),\,\textbf{v}\in H^{1}(G),$
$\left|\langle\nabla_{\Gamma}\cdot F_{\Gamma},\textbf{v}_{\Gamma}\rangle_{H^{-1}(\Gamma),H^{1}(\Gamma)}\right|\leq C_{2}\|F_{\Gamma}\|_{L^{2}(\Gamma;\mathbb{R}^{N})}\|\textbf{v}_{\Gamma}\|_{H^{1}(\Gamma)},\quad\text{for all }\;F_{\Gamma}\in L^{2}(\Gamma;\mathbb{R}^{N}),\,\textbf{v}_{\Gamma}\in H^{1}(\Gamma),$
for some positive constants $C_{1}$ and $C_{2}$.
This paper is organized as follows. In Section 2, we establish the main Carleman estimate presented in Theorem 1.1. In Section 3, we address the necessary observability problems, including the unique continuation property and the observability inequality for the adjoint equation associated with (1.7). In Section 4, we use the classical duality argument to deduce the proofs of our main controllability results i.e., Theorems 1.2 and 1.3. Finally, in Appendix A, we complete the proof of the known Carleman estimate from [6, Theorem 4.1] by extending the Laplacian operators to our general second-order diffusion operators and carefully analyzing the dependence of the parameters on the final control time $T$. This estimate plays a crucial role in Section 2.
2 Carleman Estimate for Backward Stochastic Parabolic Equations: Proof of Theorem 1.1
This section is devoted to proving the main Carleman estimate in Theorem 1.1. To begin, we first establish the following Carleman estimate for system (1.2) without divergence source terms (i.e., with $F\equiv F_{\Gamma}\equiv 0$).
@@
$\begin{aligned} &\,\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}\,dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}\,d\sigma dt\\ &+\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}\,dxdt+\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}\,d\sigma dt\\ &\leq C\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}z^{2}\,dxdt+\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}\,dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}\,d\sigma dt\\ &\qquad\quad+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}\,dxdt+\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}\,d\sigma dt\Bigg],\end{aligned}$ (2.1)
for large enough $\mu\geq C$ and $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$.
Since the proof of the above estimate (2.1) involves several technical steps, we include all details in Appendix A.
Based on the result of Lemma 2.1, we now proceed to prove Theorem 1.1. First, we establish a controllability result for the following controlled forward stochastic parabolic equation with dynamic boundary conditions and source terms:
$\begin{cases}\begin{array}[]{ll}dy-\nabla\cdot(\mathcal{A}\nabla y)\,dt=(\lambda^{3}\mu^{4}\theta^{2}\varphi^{3}z+\mathbbm{1}_{G_{0}}u)\,dt+v_{1}\,dW(t)&\text{in}\,Q,\\ dy_{\Gamma}-\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\Gamma})\,dt+\partial_{\nu}^{\mathcal{A}}y\,dt=\lambda^{3}\mu^{3}\theta^{2}\varphi^{3}z_{\Gamma}\,dt+v_{2}\,dW(t)&\text{on}\,\Sigma,\\ y_{\Gamma}=y|_{\Gamma}&\text{on}\,\Sigma,\\ (y,y_{\Gamma})|_{t=0}=(0,0)&\text{in}\,G\times\Gamma,\end{array}\end{cases}$ (2.2)
@@
There exists a triple of controls $(\widehat{u},\widehat{v}_{1},\widehat{v}_{2})\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$ such that the corresponding solution $(\widehat{y},\widehat{y}_{\Gamma})$ of (2.2) satisfies
Moreover, there exists a constant $C>0$, depending only on $G$, $G_{0}$, $\mathcal{A}$, and $\mathcal{A}_{\Gamma}$, such that the following inequality holds:
$\begin{aligned} &\,\lambda^{-3}\mu^{-4}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-3}\widehat{u}^{2}\,dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}\widehat{v}_{1}^{2}\,dxdt\\ &+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}\widehat{v}_{2}^{2}\,d\sigma dt+\mathbb{E}\iint_{Q}\theta^{-2}\widehat{y}^{2}\,dxdt+\mathbb{E}\iint_{\Sigma}\theta^{-2}\widehat{y}_{\Gamma}^{2}\,d\sigma dt\\ &+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}|\nabla\widehat{y}|^{2}\,dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}|\nabla_{\Gamma}\widehat{y}_{\Gamma}|^{2}\,d\sigma dt\\ &\leq C\bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}\,dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}\,d\sigma dt\bigg],\end{aligned}$ (2.3)
for large $\mu\geq C$ and $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$.

### v2

@@ 1 Introduction and Main Results
The surface diffusion term $\sum_{j,k=1}^{N}D_{j}\Big(a_{\Gamma}^{jk}(x)D_{k}z_{\Gamma}\Big)\,dt$” in the boundary equation of (1.2) plays a crucial role in the derivation of the Carleman estimate (1.8), in particular in Step 2 of the proof of Lemma 3.3. This term is essential for absorbing the following problematic boundary integral:
$\lambda\mu\,\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\,\bigl|\partial_{\nu}^{\mathcal{A}}\psi\bigr|\,\bigl|\nabla_{\Gamma}z_{\Gamma}\bigr|^{2}\,d\sigma\,dt.$
Consequently, the null and insensitizing control problems for (1.9) under the dynamic boundary condition
$\begin{cases}dy_{\Gamma}+\partial_{\nu}^{\mathcal{A}}y\,dt=(a_{2}y_{\Gamma}+B_{2}\cdot\nabla_{\Gamma}y_{\Gamma})\,dt+v_{2}\,dW(t),&\textnormal{on }\Sigma,\\ y_{\Gamma}=y|_{\Gamma},&\textnormal{on }\Sigma,\end{cases}$
remains an open problem. In contrast, for $N=1$, it is known that the controllability of stochastic heat equations can be addressed under this type of boundary condition; see [6]. This is due to the fact that, when $N=1$, the boundary $\Gamma$ is a zero-dimensional manifold, and hence the tangential gradient operator is trivial, i.e., $\nabla_{\Gamma}=0$.
@@
Let $\mathcal{O}\subset G$ and $\mathcal{O}_{\Gamma}^{1},\mathcal{O}_{\Gamma}^{2}\subset\Gamma$ be nonempty open subsets. The insensitizing control problem (1.11) with the energy functional
$\begin{aligned} \widetilde{\mathcal{E}}(y,y_{\Gamma}):=&\;\frac{1}{2}\,\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}}|\nabla y|^{2}\,dx\,dt+\frac{1}{2}\,\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}^{1}_{\Gamma}}|y_{\Gamma}|^{2}\,d\sigma\,dt\\ &\;+\frac{1}{2}\,\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}^{2}_{\Gamma}}|\nabla_{\Gamma}y_{\Gamma}|^{2}\,d\sigma\,dt\end{aligned}$
poses significant challenges. In particular, the Carleman estimate (1.8) is not sufficient to handle this functional. The main difficulty lies in the need to control a localized term $\mathbb{E}\int_{0}^{T}\int_{\widetilde{\mathcal{G}}}\kappa\,|z|^{2}\,dx\,dt$ using a higher-order term $\mathbb{E}\int_{0}^{T}\int_{\mathcal{G}}\kappa\,|\Delta z|^{2}\,dx\,dt,$ where $\widetilde{\mathcal{G}}\subset\mathcal{G}$ and $\kappa=\kappa(t,x)$ is a suitable weight function. A detailed discussion of the deterministic heat equation can be found in [22].
This paper is organized as follows. Section 2 presents some preliminaries and establishes the well-posedness of equation (1.2). In Section 3, we derive an intermediate Carleman estimate for equation (1.2) in the absence of weak divergence terms. Section 4 is devoted to the derivation of the main Carleman estimate stated in Theorem 1.1 for equation (1.2), where weak divergence terms are taken into account. Section 5 addresses the derivation of the required observability inequality for the adjoint backward equation associated with (CE), and the proof of the null controllability result stated in Theorem 1.2. Finally, Section 6 is devoted to the study of the insensitizing control problem (1.11) and to the proof of Theorem 1.3.
2 Preliminaries and Well-Posedness Results
In this section, we present some preliminaries that will be used throughout the paper and establish the well-posedness results for equations (1.2) (equivalently (1.3)).
We first recall the following lemma, which allows us to use integration by parts in space, both in the bulk domain $G$ and on the boundary $\Gamma$. For the proofs, we refer, for instance, to [16, 24].
@@ 2 Preliminaries and Well-Posedness Results
Let $F\in L^{2}(G;\mathbb{R}^{N})$ and $F_{\Gamma}\in L^{2}(\Gamma;T_{x}\Gamma)$. The divergence of $F$ (resp., the tangential divergence of $F_{\Gamma}$) can be extended as a continuous linear functional on $H^{1}(G)$ (resp., on $H^{1}(\Gamma)$) defined by
$\nabla\cdot F:H^{1}(G)\longrightarrow\mathbb{R},\quad v\longmapsto-\int_{G}F\cdot\nabla v\,dx+\langle F\cdot\nu,v|_{\Gamma}\rangle_{H^{-1/2}(\Gamma),\,H^{1/2}(\Gamma)},$
and
$\nabla_{\Gamma}\cdot F_{\Gamma}:H^{1}(\Gamma)\longrightarrow\mathbb{R},\quad v_{\Gamma}\longmapsto-\int_{\Gamma}F_{\Gamma}\cdot\nabla_{\Gamma}v_{\Gamma}\,d\sigma.$
Moreover, the following estimates hold:
$\left|\langle\nabla\cdot F,v\rangle_{(H^{1}(G))^{\prime},\,H^{1}(G)}\right|$ $\leq C_{1}\|F\|_{L^{2}(G;\mathbb{R}^{N})}\|v\|_{H^{1}(G)},$
$\left|\langle\nabla_{\Gamma}\cdot F_{\Gamma},v_{\Gamma}\rangle_{H^{-1}(\Gamma),\,H^{1}(\Gamma)}\right|$ $\leq C_{2}\|F_{\Gamma}\|_{L^{2}(\Gamma;\mathbb{R}^{N})}\|v_{\Gamma}\|_{H^{1}(\Gamma)},$
for all $v\in H^{1}(G)$ and $v_{\Gamma}\in H^{1}(\Gamma)$, and for some positive constants $C_{1}$ and $C_{2}$.
We now present the following Itô formula for stochastic processes in weak form (see, e.g., [34, Chapter 2]). This result will be useful for deriving energy estimates, as well as for establishing duality relations between forward and backward equations with general dynamic boundary conditions and drift terms in the $\mathbb{H}^{-1}$-Sobolev space.
@@
$(z(t),z_{\Gamma}(t))$ $=(z_{T},z_{\Gamma,T})-\int_{t}^{T}(\phi(s),\phi_{\Gamma}(s))\,ds-\int_{t}^{T}(Z(s),\widehat{Z}(s))\,dW(s),\quad\textnormal{a.s.},$
$(y(t),y_{\Gamma}(t))$ $=(y_{0},y_{\Gamma,0})+\int_{0}^{t}(\widetilde{\phi}(s),\widetilde{\phi}_{\Gamma}(s))\,ds+\int_{0}^{t}(\Psi(s),\widehat{\Psi}(s))\,dW(s),\quad\textnormal{a.s.}$
Then the following assertions hold:
  1. 1.
    For any $t\in[0,T]$, it holds almost surely that
    $\|(z(t),z_{\Gamma}(t))\|^{2}_{\mathbb{L}^{2}}$ $=\|(z(0),z_{\Gamma}(0))\|^{2}_{\mathbb{L}^{2}}+2\int_{0}^{t}\langle(\phi(s),\phi_{\Gamma}(s)),(z(s),z_{\Gamma}(s))\rangle_{\mathbb{H}^{-1},\mathbb{H}^{1}}\,ds$
    $\quad+2\int_{0}^{t}\langle(z(s),z_{\Gamma}(s)),(Z(s),\widehat{Z}(s))\rangle_{\mathbb{L}^{2}}\,dW(s)+\int_{0}^{t}\|(Z(s),\widehat{Z}(s))\|^{2}_{\mathbb{L}^{2}}\,ds.$
  2. 2.
    For all $t\in[0,T]$, it holds almost surely that
    $\langle(z(t),z_{\Gamma}(t)),(y(t),y_{\Gamma}(t))\rangle_{\mathbb{L}^{2}}$ $=\langle(z(0),z_{\Gamma}(0)),(y_{0},y_{\Gamma,0})\rangle_{\mathbb{L}^{2}}$
    $\quad+\int_{0}^{t}\langle(\phi(s),\phi_{\Gamma}(s)),(y(s),y_{\Gamma}(s))\rangle_{\mathbb{H}^{-1},\mathbb{H}^{1}}\,ds$
    $\quad+\int_{0}^{t}\langle(\widetilde{\phi}(s),\widetilde{\phi}_{\Gamma}(s)),(z(s),z_{\Gamma}(s))\rangle_{\mathbb{H}^{-1},\mathbb{H}^{1}}\,ds$
    $\quad+\int_{0}^{t}\langle(y(s),y_{\Gamma}(s)),(Z(s),\widehat{Z}(s))\rangle_{\mathbb{L}^{2}}\,dW(s)$
    $\quad+\int_{0}^{t}\langle(z(s),z_{\Gamma}(s)),(\Psi(s),\widehat{\Psi}(s))\rangle_{\mathbb{L}^{2}}\,dW(s)$
    $\quad+\int_{0}^{t}\langle Z(s),\Psi(s)\rangle_{\mathbb{L}^{2}}\,ds.$
In the above lemma, $\langle\cdot,\cdot\rangle_{\mathbb{L}^{2}}$ denotes the inner product in the space $\mathbb{L}^{2}$, while $\langle\cdot,\cdot\rangle_{\mathbb{H}^{-1},\mathbb{H}^{1}}$ denotes the duality pairing between $\mathbb{H}^{1}$ and its dual $\mathbb{H}^{-1}$, taken with respect to the pivot space $\mathbb{L}^{2}$.
We first consider the following unbounded linear operator on the product space $\mathbb{L}^{2}$:
$\textbf{A}=\begin{pmatrix}\nabla\cdot(\mathcal{A}\nabla)&0\\ -\partial_{\nu}^{\mathcal{A}}&\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma})\end{pmatrix},\qquad\mathcal{D}(\textbf{A})=\mathbb{H}^{2}.$
Following ideas similar to those in the proof of Proposition 2.1 in [36], one can show that A satisfies the following generation property.
Proposition 2.1.
The operator A is densely defined, self-adjoint, and dissipative on $\mathbb{L}^{2}$, and it generates an analytic $C_{0}$-semigroup $(S(t))_{t\geq 0}$ on $\mathbb{L}^{2}$. Moreover, we have $\mathcal{D}((-\textbf{A})^{1/2})=\mathbb{H}^{1}$.
We now give the definition of weak solution of (1.3) (equivalently (1.2)).
Definition 2.1.
A process
$(z,z_{\Gamma};Z,\widehat{Z})\in\left(L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))\bigcap L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})\right)\times L^{2}_{\mathcal{F}}(0,T;\mathbb{L}^{2})$
is said to be a weak solution of (1.3) if, for any $t\in[0,T]$ and all $(\zeta,\zeta_{\Gamma})\in\mathbb{H}^{1}$, it holds that
$\begin{aligned} &\int_{G}(z_{T}-z(t))\zeta\,dx+\int_{\Gamma}(z_{\Gamma,T}-z_{\Gamma}(t))\zeta_{\Gamma}\,d\sigma\\ &=\int_{t}^{T}\int_{G}\mathcal{A}\nabla z\cdot\nabla\zeta\,dx\,ds+\int_{t}^{T}\int_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}\zeta_{\Gamma}\,d\sigma\,ds\\ &\quad+\int_{t}^{T}\int_{G}F_{1}\zeta\,dx\,ds+\int_{t}^{T}\int_{\Gamma}F_{2}\zeta_{\Gamma}\,d\sigma\,ds\\ &\quad-\int_{t}^{T}\int_{G}F\cdot\nabla\zeta\,dx\,ds-\int_{t}^{T}\int_{\Gamma}F_{\Gamma}\cdot\nabla_{\Gamma}\zeta_{\Gamma}\,d\sigma\,ds\\ &\quad+\int_{t}^{T}\int_{G}Z\zeta\,dx\,dW(s)+\int_{t}^{T}\int_{\Gamma}\widehat{Z}\zeta_{\Gamma}\,d\sigma\,dW(s),\quad\textnormal{a.s.}\end{aligned}$ (2.1)
We first establish the following well-posedness result for equation (1.3) in the absence of divergence terms, that is, when $F=F_{\Gamma}=0$.
Proposition 2.2.
Let $(z_{T},z_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$, $F_{1}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G))$, and $F_{2}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$. Then equation (1.3) (with $F=F_{\Gamma}=0$) admits a unique weak solution
$(z,z_{\Gamma};Z,\widehat{Z})\in\Big(L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))\bigcap L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})\Big)\times L^{2}_{\mathcal{F}}(0,T;\mathbb{L}^{2}).$
Moreover, the following estimate holds:
@@
where $C>0$ is a constant depending only on $G$, $T$, $\beta_{0}$, $M$, and $M_{\Gamma}$.
Proof.
Observe that equation (1.3) can be rewritten in the following abstract backward stochastic evolution form:
$\begin{cases}d\mathbf{Z}=(-\textbf{A}\mathbf{Z}+\mathbf{F})\,dt+\widehat{\mathbf{Z}}\,dW(t),&t\in[0,T),\\ \mathbf{Z}(T)=(z_{T},z_{\Gamma,T}),\end{cases}$
[…]
We next establish the well-posedness of the full system (1.3), or equivalently, (1.2).
@@
Let $(z_{T},z_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$, $F_{1}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G))$, $F_{2}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$, $F\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G;\mathbb{R}^{N}))$, and $F_{\Gamma}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma;\mathbb{R}^{N}))$. Then equation (1.3) admits a unique weak solution
$(z,z_{\Gamma};Z,\widehat{Z})\in\Big(L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))\bigcap L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})\Big)\times L^{2}_{\mathcal{F}}(0,T;\mathbb{L}^{2}).$
Moreover, the following estimate holds:
[…]
where $C>0$ depends only on $G$, $T$, $\beta_{0}$, $M$, and $M_{\Gamma}$.